# Excitable quantum systems: the bosonic avalanche laser

### v2

@@ Abstract
We investigate the dynamics of a lasing system that is driven by a current of bosonic (quasi-) particles via a dissipative three-mode mixing process. A semi-classical analysis of this system predicts distinct dynamical regimes, where both the cavity mode and the gain medium can undergo lasing transitions. Of particular interest is an intermediate self-pulsing phase that exhibits the characteristics of an excitable system and converts random input signals into separated, quasi-periodic pulses at the output. By performing exact Monte-Carlo simulations, we extend this analysis into the quantum regime and show that despite being dominated by huge bosonic particle number fluctuations, this effect of coherence resonance survives even for rather low average photon numbers. Our system thus represents a intriguing model for an excitable quantum many-body system, with practical relevance for quantum detectors or autonomous quantum machines. As an illustration, we discuss the realization of this system with superconducting quantum circuits and its application as a number-resolved avalanche detector for microwave photons.
@@
Excitable systems are a broad class of nonlinear dynamical systems that, in simple terms, can support propagating waves or other collective excitations, but cannot be re-excited until a certain amount of time has passed. Typical physical signatures of these systems are the effects of coherence resonance (CR) and stochastic resonance (SR), whereby a purely noisy input can either generate a highly regular response or amplify a weak periodic signal Gang et al. (1993); Pikovsky and Kurths (1997); Lindner et al. (2004); Wellens et al. (2004). Excitable systems are widely studied in various areas of classical physics, in particular in the context of chemical reactions and biological systems, while comparably little is still known about such systems and their potential applications in the quantum regime. SR has been studied mainly in connection with thermally or externally driven transitions in bistable quantum systems Löfstedt and Coppersmith (1994); Grifoni and Hänggi (1996); Grifoni et al. (1996); Whitney et al. (2011); Xie et al. (2018); Hussein et al. (2020); Kato and Nakao (2021); Hänze et al. (2021); Li et al. (2024) and noise-assisted entanglement schemes Huelga and Plenio (2007); Rivas et al. (2009). CR has been observed in current oscillations in electronic lattices Mompo et al. (2018); Bonilla et al. (2024), but still in a regime where the externally applied classical noise dominates over intrinsic quantum mechanical fluctuations.
@@
In this Letter we analyze the dynamics of a novel lasing system that is driven by a dissipative current of bosonic (quasi-) particles, as shown in Fig. 1. In this system, each of the injected bosons transitions between multiple intermediate energy levels, thereby emitting multiple photons into the cavity mode before leaving the system again, similar to the concept of a bosonic cascade laser Liew et al. (2013); Kaliteevskii and Ivanov (2013); Kaliteevski et al. (2014); Kavokin et al. (2016); Liew et al. (2016). However, here we assume that each transition is assisted by a reservoir, which makes this process dissipative and, importantly, unidirectional. Under this condition, a non-trivial interplay between the bosonic current—which drives the cavity mode—and the cavity mode—which stimulates the current—emerges and gives rise to different lasing regimes. These include, in particular, a new dynamical self-pulsing phase, where within a classical mean-field description, the system emits bursts of photons with a period that is a universal function of the underlying system parameters. Each burst is followed by a dark period of no emission, as characteristic for an excitable system.
With the help of exact Monte-Carlo simulations, we study the lasing system also for moderate and low excitation numbers, where the dynamics is dominated by the intrinsic shot-noise fluctuations from the discreteness of the bosonic particle current. Even in this regime, we observe the phenomenon of CR, where those quantum fluctuations are converted into a semi-regular output signal. We further propose an implementation of this lasing system with superconducting circuits, where the described mechanism prevails even at the few-photon level and can thus find practical applications as a number-resolved avalanche detector for microwave photons.
@@
Model.—We consider the generic setup shown in Fig. 1, where a ladder of $N$ bosonic modes, the gain medium, is coupled to a lasing cavity via a dissipative three-wave mixing process. We model this system by a master equation of the form
@@
$\begin{split}\dot{\rho}=\,&\Gamma\sum_{p=1}^{N-1}\mathcal{D}[a_{p}a^{\dagger}_{p+1}c^{\dagger}]\rho+\kappa_{c}\mathcal{D}[c]\rho+\mathcal{L}_{\rm gain}\rho,\end{split}$ (1)
@@
where $\rho$ is the system density operator, $\mathcal{D}[C]\rho=C\rho C^{\dagger}-(C^{\dagger}C\rho-\rho C^{\dagger}C)/2$, and $a_{p}$ ($a^{\dagger}_{p}$) and $c$ ($c^{\dagger}$) denote the bosonic annihilation (creation) operators of the ladder modes and the cavity mode, respectively. The first term in Eq. (1) describes bosons hopping from mode $p$ to the next lower mode $p+1$, while simultaneously emitting a photon into the cavity. This process is purely incoherent and irreversible such that re-absorption events of the form $\mathcal{D}[a^{\dagger}_{p}a_{p+1}c]$ do not occur. The photons in the lasing mode decay with a rate $\kappa_{c}$ and, for a steady operation, gain bosons are injected countinuously into mode $p=1$ with rate $\gamma_{\textsl{g}}$ and leave the system through mode $p=N$ with rate $\kappa_{\ell}$. We model both processes by the Lindbladian,
$\mathcal{L}_{\rm gain}\rho=\kappa_{\ell}\mathcal{D}[a_{N}]\rho+\gamma_{\textsl{g}}(1-\zeta)\mathcal{D}[a_{1}]\rho+\gamma_{\textsl{g}}\zeta\mathcal{D}[a^{\dagger}_{1}]\rho,$ (2)
@@
where $\zeta\in[0,1]$. In the following we focus primarily on the case $\zeta=1/2$, where the gain medium is coupled to an effective infinite-temperature reservoir. However, all effects discussed below are also observed for finite temperatures $(\zeta<1/2)$ or for models with pure gain $(\zeta=1)$.
@@
Lasing regimes.—To evaluate the dynamics and steady states of this system, we start with a mean-field analysis, where we describe the cavity mode by its classical amplitude, $\alpha_{c}=\langle c\rangle$, and we neglect correlations between the populations of the different ladder modes $n_{p}=\langle a^{\dagger}_{p}a_{p}\rangle$. Under this approximation, we obtain the set of coupled equations,
@@
where $J_{p,p+1}=\Gamma n_{p}(1+n_{p+1})$ is the bosonic current between sites $p$ and $p+1$, and $J_{\rm cum}=\sum_{p}J_{p,p+1}$ the total cumulative current. Eq. (3) clearly displays double stimulation Kaliteevskii and Ivanov (2013); Kaliteevski et al. (2014), where the hopping of particles from site $p$ to $p+1$ is stimulated by both the cavity population $n_{c}=\lvert\alpha_{c}\rvert^{2}$ and the number of bosons already present on the target site $n_{p+1}$.
@@
In Fig. 2 we show the transient dynamics and steady states of the mean-field equations for $N=10$ and varying pump and decay rates. In the limit $\kappa_{c}/\gamma_{\textsl{g}}\rightarrow 0$, the cumulative current, $J_{\rm cum}$, exceeds the cavity loss rate and the lasing mode is amplified. In turn, the amplified lasing mode accelerates the current in the ladder. This interplay results in an oscillatory initial dynamics, but the system quickly relaxes into a regular, i.e., stationary, lasing state. In the opposite limit, $\kappa_{c}/\gamma_{\textsl{g}}\rightarrow\infty$, the cavity is overdamped and stays de-excited. The remaining transport dynamics of the ladder modes in Eq. (3) then describes a so-called asymmetric simple inclusion process (ASIP) Garbe et al. (2024); Minoguchi et al. (2025). Interestingly, this transport process by itself supports a lasing phase, where bosons condense into a few of the lowest modes with a characteristic zig-zag density profile Liew et al. (2016); Garbe et al. (2024). This pattern, captured by the staggered population $n_{\rm stag}=\sum_{p}(-1)^{p}(n_{p}-n_{1})$, is clearly visible in the density profile of the ladder modes shown in Fig. 2. Note that a similar staggered configuration is also found in the lasing phase Sup .
Figure 2: Mean-field phase diagram of the bosonic avalanche laser, as obtained from the solution of Eq. (3) for an initial seed amplitude $\alpha_{c}(t=0)=\sqrt{10}$. The different phases are distinguished by their characteristic dynamics, shown in the panels on the left for $\gamma_{\textsl{g}}/\Gamma=2,12,40$ and $\kappa_{c}/\Gamma=20$, and by the order parameters $\alpha_{c}$ and $n_{\rm stag}$ in the steady state. Here, $|\alpha_{c}|>0$ indicates lasing of the cavity mode and $n_{\rm stag}\neq 0$ condensation of the ladder modes Garbe et al. (2024) in a staggered density pattern, which is shown in the panel below for $\gamma_{\textsl{g}}/\Gamma=5$ and $\kappa_{c}/\Gamma=40$. In the self-pulsing phase (orange), no stationary state is reached. The other parameters used in these plots are $\kappa_{\ell}/\Gamma=10$, $\zeta=1/2$ and $N=10$.
@@
Self-pulsing regime.—In between these two limiting cases, i.e., for $\kappa_{c}\sim\gamma_{\textsl{g}}$, the system does not reach a steady state. Instead, we observe a persistent emission of periodic photon bursts, which are separated by longer gaps over which the lasing mode remains de-excited. This behavior is self-sustained and characterized by a period $\tau\equiv\tau(\kappa_{c},\gamma_{\textsl{g}},N,\Gamma)$, which is independent of the initial conditions in our mean-field simulations. Note that this behavior, as well as all the other features of the stationary phases discussed above, are robust with respect to intrinsic losses of the ladder modes with rate $\kappa_{0}\lesssim\Gamma$ Sup .
@@
To understand the origin of this behavior, we plot in Fig. 3 the individual populations $n_{p}$ over several periods. Starting from an empty system, bosons are injected into mode $p=1$, and a density wave starts to flow through the ladder with a speed $c\approx(1+n_{1})\Gamma$, enhanced by the quasi-stationary occupation $n_{1}$ of the first few modes. This number is determined by the left boundary condition, $\Gamma n_{1}(1+n_{1})=\gamma_{\textsl{g}}$. During this initial phase, $J_{\rm cum}\sim\gamma_{\textsl{g}}ct$ increases approximately linearly with time and eventually exceeds the cavity loss, $\kappa_{c}$. Beyond this point, the lasing mode is amplified; once the cavity population has become significantly larger than unity, the bosonic density wave is accelerated accordingly to $c^{\prime}\approx(1+|\alpha_{c}|^{2})c\gg c$. This speed-up empties the ladder almost instantaneously, after which the cavity mode decays, and the cycle starts anew. In view of the accelerated bosonic transport during the lasing burst, we expect the period $\tau$ to be approximately determined by the condition $J_{\rm cum}(\tau)\approx\kappa_{c}$, where $J_{\rm cum}(\tau)\approx\gamma_{\textsl{g}}(1+n_{1})\Gamma\tau$ during the initial buildup stage. For $n_{1}\gg 1$, we then obtain a scaling $\sqrt{\gamma_{\textsl{g}}\Gamma}\tau\sim(\kappa_{c}/\gamma_{\textsl{g}})$, roughly independent of the number of ladder modes. In the inset of Fig. 3, we plot the rescaled mean-field period $\sqrt{\gamma_{\textsl{g}}\Gamma}\tau$ for various different ratios between the parameters $\Gamma$, $\gamma_{\textsl{g}}$ and $\kappa_{c}$ and two different values of $N$. We observe indeed a collapse of all results onto a single universal function, although the precise dependence deviates from the estimate above, which oversimplifies the actual evolution of the gain current and the cavity mode during the build-up stage Sup .
@@
Figure 3: Origin of self-pulsing. The panel at the right bottom shows the time evolution of the mode occupation numbers $n_{p}$ over a few cycles for $\kappa_{c}/\Gamma=\kappa_{\ell}/\Gamma=20$ and $\gamma_{\textsl{g}}/\Gamma=10$. The corresponding population of the cavity mode $n_{c}$ is shown on top. In each cycle, the system evolves through three distinct phases, which are illustrated by the corresponding sketches on the left. The inset shows the rescaled pulsing period $\tau$ obtained from mean-field simulations. The various symbols represent different combination of $\kappa_{c}/\kappa_{\ell}=5,25$ and $\Gamma/\kappa_{\ell}=0.05,5$, keeping $\kappa_{\ell}$ fixed. Additional parameter combinations are shown in Sup . The colors correspond to different values of $N=10$ (red) and $N=20$ (blue). Upon rescaling, all curves collapse to the same universal behavior. For all plots, $\zeta=1/2$.

### v3

@@ Abstract
We investigate the dynamics of a lasing system driven by a current of bosonic (quasi-)particles via a dissipative three-mode mixing process. A semi-classical analysis of this system predicts distinct dynamical regimes, where both the cavity mode and the gain medium can undergo lasing transitions. Of particular interest is an intermediate self-pulsing phase that exhibits the characteristics of an excitable system and converts random input signals into separated, quasi-periodic pulses at the output. By performing exact Monte-Carlo simulations, we extend this analysis into the quantum regime and show that despite being dominated by huge bosonic particle number fluctuations, this effect—reminiscent of coherence resonance— survives even for rather low average photon numbers. Our system thus represents an intriguing model of an excitable quantum many-body system, with practical relevance for quantum detectors or autonomous quantum machines. As an illustration, we discuss the realization of this system with superconducting quantum circuits and its application as a number-resolved avalanche detector for microwave photons.
I Introduction
@@
Excitable systems are a broad class of nonlinear dynamical systems that, in simple terms, can support propagating waves or other collective excitations, but cannot be re-excited until a certain amount of time has passed. Typical physical signatures of these systems are the effects of coherence resonance (CR) and stochastic resonance (SR), whereby a purely noisy input can either generate a highly regular response or amplify a weak periodic signal [6, 27, 22, 32]. Excitable systems are widely studied in various areas of classical physics, in particular in the context of chemical reactions and biological systems; however, comparably little is still known about such systems and their potential applications in the quantum regime. SR has been studied mainly in connection with thermally or externally driven transitions in bistable quantum systems [23, 10, 11, 33, 34, 14, 17, 12, 19] and noise-assisted entanglement schemes [13, 29]. CR has been observed in current oscillations in electronic lattices [26, 2], but still in a regime where the externally applied classical noise dominates over intrinsic quantum mechanical fluctuations.
@@
In this paper, we analyze the dynamics of a novel lasing system that is driven by a dissipative current of bosonic (quasi-) particles, as shown in Fig. 1. In this system, each of the injected bosons transitions between multiple intermediate energy levels, thereby emitting multiple photons into the cavity mode before leaving the system again. This process is similar to the concept of a bosonic cascade laser [21, 16, 15, 18, 20], but here we assume that each transition is assisted by an auxiliary reservoir. This makes the transitions dissipative and, importantly, unidirectional. Under this condition, a non-trivial interplay between the bosonic current—which drives the cavity mode—and the cavity mode—which stimulates the current—emerges and gives rise to different lasing regimes. These include, in particular, a new dynamical self-pulsing phase, where, within a classical mean-field description, the system emits bursts of photons with a period that is a universal function of the underlying system parameters. Each burst is followed by a dark period with no emission, as is characteristic for excitable systems.
@@
With the help of exact Monte-Carlo simulations, we study the lasing system also for moderate and low excitation numbers, where the dynamics is dominated by the intrinsic shot-noise fluctuations that arise from the discreteness of the bosonic particle current. From these simulations, we observe that even for very low average excitation numbers, signatures of self-pulsing survive, although the pulses are strongly affected by the noise. By analyzing in more detail the correlation spectrum of the output signal and the corresponding statistics of the emission burst, we show that the behavior of this lasing system is reminiscent of the phenomenon of CR: randomly injected bosons are converted into a semi-regular output signal. Specifically, we find that the coherence parameter, a measure of the regularity of the output pulses, exhibits a maximum as a function of pumping strength. Comparing different pumping mechanisms, we further show that random fluctuations in the injection can directly improve the regularity of the output, which is a key feature of excitable systems.
Deep into the quantum regime, we find that already a single boson injected into the system can produce an avalanche of signal photons—a mechanism that can be used as an amplifier for weak quantum signals. As a specific application, we propose an implementation of the described lasing system using superconducting quantum circuits and show that, under realistic experimental conditions, this device can serve, for example, as a number-resolved detector for microwave photons.
II A bosonic avalanche laser
@@
We consider the generic setup shown in Fig. 1, where a ladder of $N$ bosonic modes, the gain medium, is coupled to a lasing cavity via a dissipative three-wave mixing process. We model this system by a master equation of the form
@@
$\begin{split}\dot{\rho}=\,&\Gamma\sum_{p=1}^{N-1}\mathcal{D}[a_{p}a^{\dagger}_{p+1}c^{\dagger}]\rho+\kappa_{c}\mathcal{D}[c]\rho+\mathcal{L}_{\rm gain}\rho+\mathcal{L}_{\rm loss}\rho,\end{split}$ (1)
@@
where $\rho$ is the system density operator, $\mathcal{D}[C]\rho=C\rho C^{\dagger}-(C^{\dagger}C\rho-\rho C^{\dagger}C)/2$, and $a_{p}$ ($a^{\dagger}_{p}$) and $c$ ($c^{\dagger}$) denote the bosonic annihilation (creation) operators of the ladder modes and the cavity mode, respectively. The first term in Eq. (1) describes bosons hopping from mode $p$ to the next lower mode $p+1$, while simultaneously emitting a photon into the cavity. This process is purely incoherent and irreversible, such that re-absorption events of the form $\mathcal{D}[a^{\dagger}_{p}a_{p+1}c]$ do not occur. In Sec. IV we discuss how, quite generically, such a dissipation process arises from a higher-order coupling to auxiliary bath modes. We also describe a more specific implementation of this process with superconducting circuits.
The photons in the cavity mode decay with a rate $\kappa_{c}$. For a steady operation, gain bosons are injected continuously into mode $p=1$ with rate $\gamma_{\textsl{g}}$ and leave the system through mode $p=N$ with rate $\kappa_{\ell}$. We model those processes by the Lindbladians
$\mathcal{L}_{\rm loss}\rho=\kappa_{\ell}\mathcal{D}[a_{N}]\rho$ (2)
for loss and
$\mathcal{L}_{\rm gain}\rho=\gamma_{\textsl{g}}(1-\zeta)\mathcal{D}[a_{1}]\rho+\gamma_{\textsl{g}}\zeta\mathcal{D}[a^{\dagger}_{1}]\rho,$ (3)
@@
for gain, respectively, where $\zeta\in[0,1]$. In the following, we focus primarily on the case $\zeta=1/2$, where the gain medium is coupled to an effective infinite-temperature reservoir. However, all effects discussed below are also observed for finite temperatures $(\zeta<1/2)$ or for models with pure gain $(\zeta=1)$.
II.1 Lasing regimes
@@
To describe the dynamics and steady state of this system, we start with a mean-field analysis, where we describe the cavity mode by its classical amplitude, $\alpha_{c}=\langle c\rangle$, and we neglect correlations between the populations of the different ladder modes $n_{p}=\langle a^{\dagger}_{p}a_{p}\rangle$. Under this approximation, we obtain the set of coupled equations,
@@
where $J_{p,p+1}=\Gamma n_{p}(1+n_{p+1})$ is the bosonic current between sites $p$ and $p+1$, $\delta_{i,j}$ is the Kronecker delta, and
$J_{\rm cum}=\sum_{p=1}^{N-1}J_{p,p+1}$ (5)
@@
is the total cumulative current that is flowing through the device. Eq. (4) clearly displays double stimulation [16, 15], where the hopping of particles from site $p$ to $p+1$ is stimulated by both the cavity population $n_{c}=\lvert\alpha_{c}\rvert^{2}$ and the number of bosons already present on the target site $n_{p+1}$.
@@
In Fig. 2 we show the transient dynamics and steady state of the mean-field equations for $N=10$ and varying pump and decay rates. In the limit $\kappa_{c}/\gamma_{\textsl{g}}\rightarrow 0$, the cumulative current, $J_{\rm cum}$, exceeds the cavity loss rate; the cavity mode is amplified and starts lasing. In turn, the amplified cavity mode accelerates the current through the ladder. This interplay results in an oscillatory initial dynamics, but the system quickly relaxes into a regular, i.e., stationary, lasing state of the cavity. In the opposite limit, $\kappa_{c}/\gamma_{\textsl{g}}\rightarrow\infty$, the cavity is overdamped and stays de-excited. The remaining transport dynamics of the ladder modes in Eq. (4) then describes a so-called asymmetric simple inclusion process (ASIP), whose transient and stationary features have been analyzed in Refs. [7, 25]. Interestingly, this transport process can itself lead to lasing within the ladder, where bosons condense into a few of the lowest modes with a characteristic zig-zag density profile [20, 7]. This pattern, captured by the staggered population $n_{\rm stag}=\sum_{p}(-1)^{p}(n_{p}-n_{1})$, is clearly visible in the density profile of the ladder modes shown in Fig. 2. Note that a similar staggered configuration is also found in the cavity lasing phase, where the cavity mode can be replaced by a $c$-number that simply enhances the rate $\Gamma$ (see Appendix A). To avoid confusion, in the rest of the manuscript, we will use the terms "lasing" and "lasing mode" only to refer to the cavity mode.
@@
In between these two limiting cases, i.e., for $\kappa_{c}\sim\gamma_{\textsl{g}}$, the system does not reach a steady state. Instead, we observe a persistent emission of periodic photon bursts, which are separated by longer gaps during which the cavity mode remains de-excited. This behavior is self-sustained and characterized by a period $\tau\equiv\tau(\kappa_{c},\gamma_{\textsl{g}},N,\Gamma)$, which is independent of the initial conditions in our mean-field simulations. Note that this behavior, as well as all the other features of the stationary phases discussed above, is robust with respect to intrinsic losses of the ladder modes, as long as those occur at a rate $\kappa_{0}\lesssim\Gamma$ (see Appendix B).
Figure 2: Mean-field phase diagram of the bosonic avalanche laser, as obtained from the solution of Eq. (4) (we used an initial seed amplitude $\alpha_{c}(t=0)=\sqrt{10}$ here, but the results are independent of the specific initial value). The different phases are distinguished by their characteristic dynamics, shown in the panels on the left for $\gamma_{\textsl{g}}/\Gamma=2,12,40$ and $\kappa_{c}/\Gamma=20$, and by the order parameters $\alpha_{c}$ and $n_{\rm stag}$ in the steady state. Here, $|\alpha_{c}|>0$ indicates lasing of the cavity mode and $n_{\rm stag}\neq 0$ condensation of the ladder modes [7] into a staggered density pattern, which is shown in the panel below for $\gamma_{\textsl{g}}/\Gamma=5$ and $\kappa_{c}/\Gamma=40$. In the self-pulsing phase (orange), no stationary state is reached. The other parameters used in these plots are $\kappa_{\ell}/\Gamma=10$, $\zeta=1/2$ and $N=10$.
II.2 Self-pulsing
@@
To understand the origin of the pulsing behavior, we plot in Fig. 3 the individual populations $n_{p}$ over several periods. The dynamics can be decomposed into three stages, sketched on the left-hand side of Fig. 3. (i) Starting from an empty system, bosons are injected into mode $p=1$ and a density wave starts to flow through the ladder with a speed $c\approx(1+\bar{n})\Gamma$, enhanced by the quasi-stationary occupation $\bar{n}$ of the first few modes. Assuming this occupation to be roughly constant throughout the packet, it can be determined by the left boundary condition, which reads $\Gamma\bar{n}(1+\bar{n})=\gamma_{\textsl{g}}$. During this initial phase, the cumulative current $J_{\rm cum}\sim\gamma_{\textsl{g}}ct$ increases approximately linearly with time, and eventually exceeds the cavity loss, $\kappa_{c}$. (ii) Beyond this point, the lasing mode is amplified; once the cavity population has become significantly larger than unity, the bosonic density wave is accelerated accordingly to a speed $c^{\prime}\approx(1+|\alpha_{c}|^{2})c\gg c$. (iii) This speed-up empties the ladder almost instantaneously, after which the cavity mode decays, and the cycle starts anew.
Given that the dynamics of the lasing burst itself is quite fast, we expect the period $\tau$ of the total process to be mainly determined by the initial build-up stage. We can express this condition approximately as $J_{\rm cum}(\tau)\approx\kappa_{c}$, with $J_{\rm cum}(\tau)\approx\gamma_{\textsl{g}}(1+\bar{n})\Gamma\tau$. For $\bar{n}\gg 1$, we then obtain a scaling $\sqrt{\gamma_{\textsl{g}}\Gamma}\tau\sim(\kappa_{c}/\gamma_{\textsl{g}})$, roughly independent of the number of ladder modes. In the inset of Fig. 3, we plot the rescaled mean-field period $\sqrt{\gamma_{\textsl{g}}\Gamma}\tau$ for various different ratios between the parameters $\Gamma$, $\gamma_{\textsl{g}}$, and $\kappa_{c}$, and two different values of $N$. We observe indeed a collapse of all results onto a single universal function, although the precise dependence deviates from the estimate above, which oversimplifies the actual evolution of the gain current and the cavity mode during the build-up stage. A more detailed description of the pulse-generating mechanism, and of the scaling of the pulsing period, can be found in Appendix C.
@@
Figure 3: Origin of self-pulsing. The panel on the bottom right shows the time evolution of the mode occupation numbers $n_{p}$ over a few cycles for $\kappa_{c}/\Gamma=\kappa_{\ell}/\Gamma=20$ and $\gamma_{\textsl{g}}/\Gamma=10$. The corresponding population of the cavity mode, $n_{c}$, is shown on top. In each cycle, the system evolves through three distinct phases, which are illustrated by the corresponding sketches on the left. The inset shows the rescaled pulsing period $\tau$ obtained from mean-field simulations. The various symbols represent different combinations of $\kappa_{c}/\kappa_{\ell}=5,25$ and $\Gamma/\kappa_{\ell}=0.05,5$, keeping $\kappa_{\ell}$ fixed. Additional parameter combinations are shown in Appendix C. The colors correspond to different values of $N=10$ (red) and $N=20$ (blue). Upon rescaling, all curves collapse to the same universal behavior. For all plots, $\zeta=1/2$.
III Quantum stochastic dynamics